• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name danez

    Log entry time 14:53:51 on February 22, 2011

    Entry number 348106

    keyword=RHRS Scalers Update

    There were two problems with the scalers in the right arm that were
    examined and corrected.
    1) All the scalers in the S1 region were giving high rates, even the ones
    not connected to anything and the ones connected to PMTs whose HV were
    turned off.
    Upon testing, it was determined that the ribbon cable connecting the S1
    signals to the scalers were bad, and when the cable was replaced, the
    rates looked normal and reasonable, and the ones not connected went to 0.
    2) The scaler signals for S2mL were abnormally high, especially when
    compared to S2mR, which were more reasonable.
    Upon testing, it was determined that the problem was the scaler module
    itself. When the ribbon cable connected S2mL to the scalers was plugged
    into a different module, the rates returned to normal.
    Specifically, S2mL was originally plugged into "slot" 2, the first of the
    three 1151s, and was moved to "slot" 1, the SiS the luminosity scalers
    are plugged into.
    Additionally, the scaler map was updated to include these changes, so
    S2mL now shares the tab with the luminosity scalers.

    A copy of this log entry has been emailed to: yez, vasulk